Simone is now 50 years old and plan to retire at age 67 (in 17 years). She currently has a share portfolio worth $750,000, a superannuation fund worth $1,200,000, and a money market (similar to cash) account worth $500,000. Her share portfolio is expected to provide annual returns of 12% p.a. (compounded annually), her superannuation will earn her 9% p.a. (compounded annually), and the money market account earns 1.2% p.a. (compounded monthly). Assume all these returns are aftertax. Assume Simone’s superannuation contribution is $25,000 (after-tax) per year for the next 17 years (starting 1 year from now).

a. If Simone stick to her current investment choices, how much will the total value of her investments be worth when she retires at 67 (in 17 years)?

Need complete formula and calculation process.
Answer #1
FUTURE VALUE OF SHARE PORTFOLIO:
Rate Annual Return Rate 12%
Nper Number of years 17
Pv Current Value $750,000
FV1 Future Value at 67 $5,149,531 (Using FV function of excel with Rate=12%, Nper =17, Pv=-750000)
FUTURE VALUE OF MONEY MARKET ACCOUNT:
Rate Annual Return Rate 1%
Nper Number of years 17
Pv Current Value $500,000
FV2 Future Value at 67 $612,405 (Using FV function of excel with Rate=1.2%, Nper =17, Pv=-500000)
FUTURE VALUE OF SUPERANNUATION FUND:
Rate Annual Return Rate 9%
Nper Number of years 17
Pv Current Value $1,200,000
Pmt Annual Contribution $25,000
FV3 Future Value at 67 $6,117,503 (Using FV function of excel with Rate=9%, Nper =17, Pmt=-25000Pv=-120000000)
Total Value of her investment at retirement (at age 67)
FV1+FV2+FV3 $11,879,438

fe =FV(D3,D4,-D5) D6 Н FUTURE VALUE OF SHARE PORTFOLIO: Annual Return Rate 12% Rate Number of years 4 Nper 17 $750,000 $5,149Angment Mumpeт эулез Cens Cipooaru fe =FV(D9,D10,-D11) D12 Н FUTURE VALUE OF SHARE PORTFOLIO: Annual Return Rate 12% Rate Num=FV(D15,D16,-D18,-D17) D19 Н K L. FUTURE VALUE OF SHARE PORTFOLIO: Annual Return Rate 12% Rate Number of years 4 Nper 17 Curr